# VoiceBench: Benchmarking LLM-Based Voice Assistants

### v2

@@ 3.1 Dataset Construction
To evaluate the general knowledge of voice assistants, we incorporate three datasets: AlpacaEval, CommonEval, and SD-QA. For AlpacaEval (Li et al., 2023), we follow LLaMA-Omni (Fang et al., 2024) to remove questions related to mathematics and coding. For CommonEval, we manually collect information-seeking questions from CommonVoice (Ardila et al., 2020), which features speech recorded in realistic settings by a diverse range of speakers using their personal devices. SD-QA (Faisal et al., 2021) consists of spoken questions with varying accents from the TyDi-QA dataset (Clark et al., 2020). In its original format, models are required to answer these questions using context passages. However, in voice interaction scenarios, providing extensive context is often impractical. Therefore, we feed only the spoken questions to voice assistants, requiring them to respond using their internal knowledge. Upon reviewing the dataset, we observe that some questions, such as "How many people live in Dallas?", require up-to-date information that frequently changes and cannot be answered without additional context. To ensure fair evaluation, we retain only those questions that could be answered using the internal knowledge of large language models (LLMs). Specifically, we prompt three advanced LLMs—Claude-3.5 (Anthropic, 2024)222claude-3-5-sonnet@20240620, Gemini-1.5-Pro (Reid et al., 2024)333gemini-1.5-pro-002, and GPT-4o (Achiam et al., 2023)444gpt-4o-2024-08-06—to answer these questions, and we manually select those questions that received correct answers from at least one of the models.
@@
Due to the high cost of generating real spoken instructions, relying solely on actual speech to evaluate voice assistants comprehensively is challenging. As a result, after preparing the text-based instructions, we convert them into speech using text-to-speech (TTS) models. Some text elements, such as special characters, are difficult to pronounce when directly converted to speech. Therefore, we first transform the text into a spoken-friendly format. Given the strong performance of LLMs in TTS text normalization (Zhang et al., 2024), we use GPT-4o for this task. The normalization prompt used can be found in APPX. A. Once the text is normalized, we feed it into TTS models for speech generation. We use the Google TTS API [^5] https://cloud.google.com/speech-to-text for this purpose, as it is not used in the training of any of the models we are evaluating, and it produces high-quality synthetic speech in comparison to other TTS systems. Notably, our results show that voice assistants exhibit similar performance rankings across data generated by different TTS models, supporting the reliability of conducting evaluation with Google TTS synthetic data. A detailed comparison of model performances across synthetic data from various TTS systems is provided in APPX. C.
@@ 3.2 Experiment Setup
We evaluate various end-to-end voice assistants on the proposed VoiceBench, including Qwen2-Audio (Chu et al., 2024), LLaMA-Omni (Fang et al., 2024), Mini-Omni (Xie and Wu, 2024a), Mini-Omni2 (Xie and Wu, 2024b), VITA (Fu et al., 2024), and DiVA (Held et al., 2024). Additionally, we build a naive voice assistant pipeline, where an automatic speech recognizer transcribes the input speech query into text, and a text-only LLM generates a response based on the transcribed query. The architectures of the evaluated models are summarized in Tab. 2.
@@
AlpacaEval
CommonEval
SD-QA
(Panda/GPTEval)
IFEval
(Prompt/Instruct Acc.)
AdvBench
(Refusal Rate)
Overall
Text Speech Text Speech Text Speech Text Speech Text Speech Text Speech
Naive 4.69 4.53 4.38 4.04 77.76 75.41 72.33 68.54 73.91 79.52 65.37 73.70 96.54 98.08 86.25 81.88
DiVA 4.68 3.67 4.29 3.54 78.30 74.50 62.39 51.72 68.70 76.31 34.93 43.38 99.23 98.27 85.51 67.73
LLaMA-Omni 4.39 3.70 4.32 3.46 55.33 60.40 40.14 39.24 45.38 56.53 10.15 19.58 98.46 11.35 76.29 41.83
Mini-Omni 2.34 1.95 2.55 2.02 26.04 7.23 23.69 4.16 13.04 22.89 8.99 18.17 86.35 37.12 43.75 28.80
Mini-Omni2 2.65 2.32 2.86 2.18 13.02 9.76 11.03 7.59 10.15 17.87 7.25 15.86 92.88 57.50 45.70 33.67
Qwen2-Audio 4.11 3.74 3.77 3.43 61.66 40.69 41.77 29.66 28.70 38.06 20.73 31.93 96.73 96.73 67.78 60.45
VITA 4.00 3.38 3.88 2.15 72.69 76.13 31.28 24.59 48.99 57.53 18.12 27.51 95.19 26.73 76.08 37.62
Table 3: The performance of various voice assistants on VoiceBench. The Text and Speech columns refer to the model performance with text-form and speech-form instructions respectively. We report the performance on SD-QA United States accent above.
@@
Since the focus of this work is to assess the quality of output content, and not all voice assistants support speech output, we directly assess the quality of text responses instead of speech output or speech transcription quality to ensure a fair comparison. For both AlpacaEval and CommonEval open-ended question answering tasks, we use GPT to assign a score between 1 and 5 to the generated responses based on the ground-truth instructions. For SD-QA, where human-labeled reference answers are available, we calculate the accuracy of the generated responses. To determine the correctness of the SD-QA responses, we employ two methods: PANDA (Li et al., 2024b), and automatic GPT evaluation, both demonstrating a strong correlation with human judgments. For IFEval, we follow the original rule-based implementation (Zhou et al., 2023) to calculate both loose and strict accuracy, reporting the average of the two accuracies at the prompt and instruction level. For AdvBench, we use the refusal rate as a measure of the safety of voice assistants, with a higher refusal rate indicating a safer assistant. Following previous LLM safety literature (Xu et al., 2024; Zou et al., 2023), we determine refusal status based on the presence of predefined refusal phrases (e.g., "Sorry, I cannot…") in the generated responses. For all evaluations based on GPT, we utilize GPT-4o-mini666gpt-4o-mini-2024-07-18. Detailed GPT evaluation prompts can be found in APPX. B. To understand the gap between text and speech processing capabilities of voice assistants, we consider two settings: one where assistants generate responses based on ground-truth text-form instructions, and another where they respond to speech-form instructions.
@@ 3.3 Results
Firstly, the naive pipeline-based voice assistant significantly outperforms all end-to-end models on spoken instructions, with a margin exceeding 10 points. Among the end-to-end models, Mini-Omni and Mini-Omni2 underperform significantly due to its use of a smaller speech encoder and base LLM, which prioritize processing efficiency over performance. Notably, while some end-to-end models, such as Qwen2-Audio (Chu et al., 2024), demonstrate comparable or even superior ASR and audio analysis performance Yang et al. (2024) compared to models such as Whisper (Radford et al., 2023) or the naive pipeline, a significant performance gap persists when handling spoken instructions in our AudioBench evaluation. This highlights a major limitation in current voice assistant evaluations, which rely heavily on ASR metrics.
Secondly, inadequate training of voice assistants can greatly impair the text processing capabilities of large language models (LLMs). For example, LLaMA-Omni and the Naive baseline both utilize the same base LLM, yet LLaMA-Omni exhibits a significant performance drop of over 11 points across all text processing tasks after additional tuning. This performance degradation is particularly severe on instruction-following tasks, which aligns with previous research indicating that end-to-end audio LLMs often struggle with instruction following Yang et al. (2024); Chen et al. (2024b).
Thirdly, we observe a notable disparity between the text and speech processing abilities of current models. The naive pipeline model shows a relatively small performance gap of 4.37 points from text to speech instructions, primarily due to ASR errors introduced by the speech recognition sub-model within the pipeline. In contrast, end-to-end models exhibit much larger gaps. For instance, VITA demonstrates a performance gap exceeding 35 points when taking text and speech inputs.
@@
Fourthly, we identify potential safety concerns with some voice assistants in voice interaction mode. While all assistants exhibit robust behavior when handling malicious instructions in text form, several models, such as Mini-Omni, fail to reject malicious instructions when they are delivered in speech form, responding directly instead.
Finally, to examine the reliability of using synthetic data to perform evaluation, we list the performance of voice assistants on real and synthetic CommonEval in Tab. 4. The performance rank on synthetic data and real data shows a strong correlation, while there still exists some discrepancy. Notably, all models achieve better performance on synthetic data. In particular, VITA has around 50% performance improvements when switching to synthetic data. Since CommonEval speeches are recorded with personal devices, instead of professional studio devices. The real speeches are usually more noisy than synthetic data, which is more challenging. These findings suggest that VITA may struggle to perform effectively during real user interactions, reinforcing our motivation to simulate real-world speech variations to thoroughly evaluate voice assistants.
@@ Appendix C Evaluation results with different TTS models
The evaluation results on synthetic speeches by different TTS models are summarized in Tab. 7. Overall, the voice assistants achieves the best performance on data generated by Google TTS, illustrating the high quality of the Google TTS. Additionally, we observe a similar model rank across different synthetic data, which supports the reliability and validity of using Google TTS synthetic data to evaluate the performance of voice assistants.

### v3

@@ 3.1 Dataset Construction
To evaluate the general knowledge of voice assistants, we incorporate three different format of questions: open-ended QA, reference-based QA, and multiple-choice QA. For open-ended QA, we include two datasets: AlpacaEval and CommonEval. For AlpacaEval (Li et al., 2023), we follow LLaMA-Omni (Fang et al., 2024) to remove questions related to mathematics and coding. For CommonEval, we manually collect information-seeking questions from CommonVoice (Ardila et al., 2020), which features speech recorded in realistic settings by a diverse range of speakers using their personal devices. Different from open-ended QA, for reference-based QA, a human-written reference answer is available to evaluate the accuracy of model responses. We select SD-QA (Faisal et al., 2021), which consists of spoken questions with varying accents from the TyDi-QA dataset (Clark et al., 2020). In its original format, models are required to answer these questions using context passages. However, in voice interaction scenarios, providing extensive context is often impractical. Therefore, we feed only the spoken questions to voice assistants, requiring them to respond using their internal knowledge. Upon reviewing the dataset, we observe that some questions, such as "How many people live in Dallas?", require up-to-date information that frequently changes and cannot be answered without additional context. To ensure fair evaluation, we retain only those questions that could be answered using the internal knowledge of large language models (LLMs). Specifically, we prompt three advanced LLMs—Claude-3.5 (Anthropic, 2024)222claude-3-5-sonnet@20240620, Gemini-1.5-Pro (Reid et al., 2024)333gemini-1.5-pro-002, and GPT-4o (Achiam et al., 2023)444gpt-4o-2024-08-06—to answer these questions, and we manually select those questions that received correct answers from at least one of the models. For multiple-choice QA, we create two datasets: OpenBookQA and MMSU, derived from MMLU-Pro. For each sample, we concatenate the question text, the options, and an instruction: "What is the answer to the above multiple-choice question? Select one of the following: A, B, C, or D." as the model input. Given that many audio assistants struggle to handle lengthy audio inputs, we randomly sample four options to generate new options for MMSU. Additionally, we exclude instructions from the Math and Computer Science categories, as these are less suitable for voice-based interactions.
@@
Due to the high cost of generating real spoken instructions, relying solely on actual speech to evaluate voice assistants comprehensively is challenging. As a result, after preparing the text-based instructions, we convert them into speech using text-to-speech (TTS) models. Some text elements, such as special characters, are difficult to pronounce when directly converted to speech. Therefore, we first transform the text into a spoken-friendly format. Given the strong performance of LLMs in TTS text normalization (Zhang et al., 2024), we use GPT-4o for this task. The normalization prompt used can be found in APPX. A. Once the text is normalized, we feed it into TTS models for speech generation. We use the Google TTS API [^5] https://cloud.google.com/speech-to-text for this purpose, as it is not used in the training of any of the models we are evaluating, and it produces high-quality synthetic speech in comparison to other TTS systems. Notably, our results show that voice assistants exhibit similar performance rankings across data generated by different TTS models, supporting the reliability of conducting evaluation with Google TTS synthetic data. A detailed comparison of model performances across synthetic data from various TTS systems is provided in APPX. C. Finally, we manually validate the correctness of the generated spoken instructions, rewriting text instructions as needed. To ensure compatibility with the audio processing limitations of various voice assistants (Chu et al., 2024; Held et al., 2024), we cap the duration of spoken instructions in VoiceBench at 30 seconds.
@@ 3.2 Experiment Setup
We evaluate various end-to-end voice assistants on the proposed VoiceBench, including Qwen2-Audio (Chu et al., 2024), LLaMA-Omni (Fang et al., 2024), Mini-Omni (Xie and Wu, 2024a), Mini-Omni2 (Xie and Wu, 2024b), VITA (Fu et al., 2024), Moshi (Défossez et al., 2024), and DiVA (Held et al., 2024). Additionally, we build two naive voice assistant pipelines, where an automatic speech recognizer transcribes the input speech query into text, and a text-only LLM generates a response based on the transcribed query. Finally, we include proprietary GPT-4o-Audio in the evaluation. The architectures of the evaluated voice assistants are summarized in Tab. 2.
@@
Model AlpacaEval CommonEval SD-QA MMSU OpenBookQA IFEval AdvBench Overall
(GPT) (GPT) (Panda/GPT) (Acc.) (Acc) (P./I. Acc) (Refusal Rate)
Naive T. 4.69 4.38 77.76 75.41 66.23 72.53 73.91 79.52 96.54 81.43
S. 4.53 4.04 72.33 68.54 62.43 81.54 65.37 73.70 98.08 79.06
Naive-4o T. 4.83 4.63 63.47 94.39 85.17 94.29 77.68 83.43 98.27 89.49
S. 4.80 4.47 60.58 90.96 81.69 92.97 73.19 79.82 98.27 87.23
DiVA T. 4.68 4.29 78.30 74.50 63.31 76.70 68.70 76.31 99.23 81.08
S. 3.67 3.54 62.39 51.72 25.76 25.49 34.93 43.38 98.27 55.70
LLaMA-Omni T. 4.39 4.32 55.33 60.40 59.01 79.34 45.38 56.53 98.46 74.26
S. 3.70 3.46 40.14 39.24 25.93 27.47 10.15 19.58 11.35 37.51
Mini-Omni T. 2.34 2.55 26.04 7.23 26.74 30.55 13.04 22.89 86.35 39.43
S. 1.95 2.02 23.69 4.16 24.69 26.59 8.99 18.17 37.12 27.90
Mini-Omni2 T. 2.65 2.86 13.02 9.76 27.13 32.09 10.15 17.87 92.88 41.10
S. 2.32 2.18 11.03 7.59 24.27 26.59 7.25 15.86 57.50 31.32
Qwen2-Audio T. 4.11 3.77 61.66 40.69 45.02 67.91 28.70 38.06 96.73 64.55
S. 3.74 3.43 41.77 29.66 35.72 49.45 20.73 31.93 96.73 55.35
VITA T. 4.00 3.88 72.69 76.13 64.54 83.08 48.99 57.53 95.19 75.43
S. 3.38 2.15 31.28 24.59 25.70 29.01 18.12 27.51 26.73 34.68
Moshi S. 2.01 1.60 15.01 16.27 24.04 26.15 6.38 13.76 44.23 27.47
GPT-4o-Audio S. 4.78 4.49 61.12 89.87 80.25 89.23 74.86 77.17 98.65 86.42
Table 3: The performance of various voice assistants on VoiceBench. The T. and S. rows refer to the model performance with text-form and speech-form instructions respectively. GPT-4o-Audio and Moshi only allows speech-form instructions now. We report the performance on SD-QA United States accent above.
@@
Since the focus of this work is to assess the quality of output content, and not all voice assistants support speech output, we directly assess the quality of text responses instead of speech output or speech transcription quality to ensure a fair comparison. For both AlpacaEval and CommonEval open-ended question answering tasks, we use GPT to assign a score between 1 and 5 to the generated responses based on the ground-truth instructions. For SD-QA, where human-labeled reference answers are available, we calculate the accuracy of the generated responses. To determine the correctness of the SD-QA responses, we employ two methods: PANDA (Li et al., 2024b), and automatic GPT evaluation, both demonstrating a strong correlation with human judgments. For OpenBookQA and MMSU, we use a rule-based method to extract the answer option (i.e., A, B, C, or D) from the model’s responses and calculate the accuracy based on these extracted answers. For IFEval, we follow the original rule-based implementation (Zhou et al., 2023) to calculate both loose and strict accuracy, reporting the average of the two accuracies at the prompt and instruction level. For AdvBench, we use the refusal rate as a measure of the safety of voice assistants, with a higher refusal rate indicating a safer assistant. Following previous LLM safety literature (Xu et al., 2024; Zou et al., 2023), we determine refusal status based on the presence of predefined refusal phrases (e.g., "Sorry, I cannot…") in the generated responses. For all evaluations based on GPT, we utilize GPT-4o-mini666gpt-4o-mini-2024-07-18. Detailed GPT evaluation prompts can be found in APPX. B. To understand the gap between text and speech processing capabilities of voice assistants, we consider two settings: one where assistants generate responses based on ground-truth text-form instructions, and another where they respond to speech-form instructions.
@@ 3.3 Results
Pipelines outperform E2E models:
Both naive pipeline-based voice assistant significantly outperforms all open-source end-to-end models on spoken instructions, with a large margin exceeding 20 points. The proprietary GPT-4o-Audio, although it still lags slightly behind its pipeline counterpart Naive-4o, achieves an exceptionally small performance gap, showcasing its superiority over existing open-source voice assistants. Among the open-source end-to-end models, Mini-Omni and Mini-Omni2 underperform significantly due to its use of a smaller speech encoder and base LLM, which prioritize processing efficiency over performance. Notably, while some end-to-end models, such as Qwen2-Audio (Chu et al., 2024), demonstrate comparable or even superior ASR and audio analysis performance Yang et al. (2024) compared to models such as Whisper (Radford et al., 2023) or the naive pipeline, a significant performance gap persists when handling spoken instructions in our AudioBench evaluation. This highlights a major limitation in current voice assistant evaluations, which rely heavily on ASR metrics.
Training impacts text instruction performance:
Inadequate training of voice assistants can greatly impair the text processing capabilities of LLMs. For example, LLaMA-Omni and the Naive baseline both utilize the same base LLM, yet LLaMA-Omni exhibits a significant performance drop of over 11 points across all text processing tasks after additional tuning. This performance degradation is particularly severe on instruction-following tasks, which aligns with previous research indicating that end-to-end audio LLMs often struggle with instruction following Yang et al. (2024); Chen et al. (2024b).
Text-speech performance gap:
We observe a notable disparity between the text and speech processing abilities of current end-to-end models. The naive pipeline model shows a relatively small performance gap of 4.37 points from text to speech instructions, primarily due to ASR errors introduced by the speech recognition sub-model within the pipeline. Additionally, Naive-4o demonstrates an even smaller performance gap, indicating that stronger backend LLMs exhibit greater resilience to potential ASR errors. In contrast, end-to-end models such as VITA exhibit a performance gap exceeding 35 points when handling both text and speech inputs. This gap is particularly pronounced in multiple-choice QA tasks, where most end-to-end models, except for Qwen2-Audio, perform at a level comparable to random guessing.
@@
Unsafe E2E models:
We identify potential safety concerns with some voice assistants in voice interaction mode. While all assistants exhibit robust behavior when handling malicious instructions in text form, several models, such as Mini-Omni, fail to reject malicious instructions when they are delivered in speech form, responding directly instead.
Comparison of synthetic and real instructions:
To examine the reliability of using synthetic data to perform evaluation, we list the performance of voice assistants on real and synthetic CommonEval in Tab. 4. The performance rank on synthetic data and real data shows a strong correlation, while there still exists some discrepancy. Notably, all models achieve better performance on synthetic data. In particular, VITA has around 50% performance improvements when switching to synthetic data. Since CommonEval speeches are recorded with personal devices, instead of professional studio devices. The real speeches are usually more noisy than synthetic data, which is more challenging. These findings suggest that VITA may struggle to perform effectively during real user interactions, reinforcing our motivation to simulate real-world speech variations to thoroughly evaluate voice assistants.
@@ Appendix C Evaluation results with different TTS models
The evaluation results on synthetic speeches generated by various TTS models are summarized in Tab. 7. Overall, the voice assistants demonstrate the best performance on data produced by Google TTS, highlighting the superior quality of Google’s text-to-speech system. This outcome reinforces the effectiveness of Google TTS in generating realistic and high-quality synthetic speech. Furthermore, we observe a consistent ranking of models across different synthetic datasets, which underscores the reliability and validity of using Google TTS synthetic data as a benchmark for evaluating the performance of voice assistants.